OpenStudy (anonymous):

Help :( 1. Complete the table below to solve the equation 2.5x – 10.5 = 64(0.5x). x f(x) = 2.5x – 10.5 g(x) = 64(0.5x) 2 3 4 5 6

OpenStudy (campbell_st):

well just substitute x = 1 into both equations and you get f(2) = 2.5*2 - 10.5 = -5.5 g(2) = 64(0.5*2) = 64 f(3) = 2.5*3 - 10.5 = -3 g(3) = 64(0.5*3) = 96 just repeat the process for x = 4, 5, and 6 you may see the pattern developing.

OpenStudy (anonymous):

* 64(0.5)^x